# Attitudes Towards and Knowledge of Non-Consensual Synthetic Intimate Imagery in 10 Countries

### 26 Jan 2024 v1

@@ 4.2. Behaviors
Behaviors were asked about in three different ways. First, respondents who reported that they were quite familiar with the term ”deepfakes” or knew a little bit about it (28.1%, $95\%CI=27.0\%-29.3\%$) were asked a follow-up question about whether they had engaged in specific deepfake porn behaviors. Second, as part of the larger survey, respondents were asked about general image-based abuse victimization and perpetration. If they reported experiences with digitally altered content, they were asked follow-up questions about whether the digitally altered content were videos, and if so, if the videos were deepfakes. This is distinct from the first question in that 1) it asks for more details of the perpetration/victimization (e.g., whether they had created, threatened to post, or posted such content), and 2) it asks respondents who reported victimization to identify whether they had been the victim of NSII specifically.
4.2.1. Behaviors of ”Familiar” Respondents
The results of the first question (measuring perpetration behaviors) can be seen in Figure 2, where the reported data represents the percentage of all respondents. The respondents who did not see this question due to a lack of familiarity with deepfake pornography are represented in the missing data, which would make each bar sum to 100% if included. For example, 70% of respondents in Spain indicated they had never heard of this until now or they had only heard the phrase, so they were not asked this question. The most common behavior was the viewing of celebrity deepfake pornography, ranging from 2.8% ($95\%CI=2.1\%-3.7\%$) of all respondents in Poland, to 8.0% ($95\%CI=6.6\%-9.5\%$) of all respondents in Australia. Aggregated across the different countries, 6.2% ($95\%CI=5.6\%-6.9\%$) of all respondents indicated that they had viewed celebrity deepfake pornography. The second most common behavior (2.9% of all respondents, $95\%CI=2.5\%-3.3\%$) was the viewing of deepfake pornography involving ”ordinary” people, ranging from 0.7% ($95\%CI=0.4\%-1.3\%$) in Denmark, to 5.1% ($95\%CI=4.1\%-6.2\%$) in Mexico. Behaviors that required more time, skills, or resources were rarer, such as creating or paying someone to create deepfake pornography of ordinary people (0.6%, $95\%CI=0.4\%-0.9\%$, and 0.5%, $95\%CI=0.3\%-0.7\%$), respectively, aggregated across countries)).
Figure 2. Deepfake Porn Behaviors by Country. Percentage of all respondents who indicated engaging, or not, in a particular behavior. 95% confidence interval bars presented in grey
@@ 4.2.1. Behaviors of ”Familiar” Respondents
Figure 2. Deepfake Porn Behaviors by Country. Percentage of all respondents who indicated engaging, or not, in a particular behavior. 95% confidence interval bars presented in grey
@@ 4.3.1. Victimization
Experiences broken out by country are available in Figures 4 and 3. Aggregated across behaviors, respondents from Australia (3.7%), Mexico (2.9%), and South Korea (3.1%) reported the highest rates of victimization by deepfake pornography. Full results with 95%CIs can be seen in Table A2 in the Appendix.
Aggregated across countries, 2.2% ($95\%CI=1.8\%-2.6\%$) of respondents reported some form of victimization. This included: 1.2% ($95\%CI=0.9\%-1.5\%$) of respondents who reported that someone had created NSII content of them; 1.3% ($95\%CI=1.0\%-1.6\%$) who reported that someone had posted or sent NSII content of them; and 1.2% ($95\%CI=0.9\%-1.5\%$) who reported that someone had threatened to post or send NSII content of them.
Figure 3. % of Respondents Reporting Different Types of Deepfake Video Victimization, by Country. 95% CIs in grey.
@@
Figure 3. % of Respondents Reporting Different Types of Deepfake Video Victimization, by Country. 95% CIs in grey.
@@ 4.3.2. Perpetration
Aggregating across behavior types (creating, posting, threatening to post), self-reported NSII perpetration was rare (1.8%, $95\%CI=1.4\%-2.2\%$), with the highest rates reported in Australia (2.4%, $95\%CI=1.7\%-3.4\%$) and the United States (2.6%, $95\%CI=1.9\%-3.5\%$), and lowest in Belgium (0.3%, $95\%CI=-0.1\%-0.7\%$). Full results can be seen in Table A2 in the Appendix. Aggregating across countries, 1.0% ($95\%CI=0.8\%-1.4\%$) of respondents indicated creating deepfake pornography, 1.0% ($95\%CI=0.8\%-1.3\%$) reported threatening to post, send, or share deepfake pornography, and 0.7% ($95\%CI=0.5\%-0.9\%$) reported actually posting, sending, or sharing deepfake pornography content. We included the creation of deepfake pornography as a perpetration behavior to mirror the fact that ”creation” behaviors, such as filming or photographing real content without consent, fall under the umbrella of image-based abuse perpetration. Nevertheless, we acknowledge that such behaviors might not be considered as ”perpetration” by others.
Figure 4. % of Respondents Reporting Different Types of Deepfake Video Perpetration, by Country. 95% CIs in grey.
@@
Figure 4. % of Respondents Reporting Different Types of Deepfake Video Perpetration, by Country. 95% CIs in grey.
@@ 4.4.1. Awareness and Attitudes
[…]
Figure 5. Heatmap of Mean Criminalization Attitudes by Respondent Gender. Scale ranges from -2 (”Definitely should not be a crime”) to 2 (”Definitely should be a crime”), with a midpoint of 0 (”Not sure”). Brighter red is more deserving of criminalization, darker blue is less deserving of criminalization. Standard errors are presented in parantheses.
Heatmap showing attitudes towards various image-based sexual abuse behaviors. The x axis is gender (women and men), and the y axis is each type of behavior (e.g., posting, sending, or showing real intimate content without consent). The behaviors involving real content all fall between ”should be a crime” and ”definitely should be a crime,” and all weighted means are ¿0, indicating that all of the behaviors fell between ”not sure” and ”definitely should be a crime.”
Figure 5. Heatmap of Mean Criminalization Attitudes by Respondent Gender. Scale ranges from -2 (”Definitely should not be a crime”) to 2 (”Definitely should be a crime”), with a midpoint of 0 (”Not sure”). Brighter red is more deserving of criminalization, darker blue is less deserving of criminalization. Standard errors are presented in parantheses.
@@ 4.4.2. Behaviors
Of the two most commonly reported behaviors, men respondents were more likely to report viewing deepfake pornography, both that of celebrities ($F=71.1,ndf=1,ddf=16510,p<0.001,RR=3.5,95\%CI=2.77-4.42$) and non-celebrities ($F=14.3,ndf=1,ddf=16510,p<0.001,RR=2.66,95\%CI=1.91-3.70$).
Figure 6. Deepfake Porn Behaviors by Gender. Percentage of all respondents who indicated engaging, or not, in a particular behavior. 95% confidence interval bars presented in grey.
@@
Figure 6. Deepfake Porn Behaviors by Gender. Percentage of all respondents who indicated engaging, or not, in a particular behavior. 95% confidence interval bars presented in grey.
@@ 4.4.3. Respondents who Reported Personal Victimization or Perpetration
Looking at the effects of gender on victimization and perpetration, men reported significantly higher rates across all victimization and perpetration experiences related to creation and threats. For example, the risk of men reporting having been threatened with deepfake dissemination was 2.6 times that of women, and the risk of them threatening someone else was 1.9 times that of women. Full results are presented below in Table 4.4.3. Notably, there were no significant differences by gender in the risk of disseminating content or having their content disseminated.
@@ 4.5. Qualitative Responses
In the survey, respondents were asked an optional, open-ended question: ”Would you like to add any additional thoughts on the topic of ‘deepfakes’? Please feel free to skip this question if you don’t have anything to add.” A total of 2,477 open-ended responses were collected. Of those responses, 1,180 were removed due to inappropriate or irrelevant responses. This left us with 1,296 open-ended responses (7.8% of respondents). No respondents discussed their own experiences of either victimization or perpetration.
@@
The second most prevalent theme, found in just over 30% of all responses to the open-ended question, referred to a fear of technology; specifically deepfakes and artificial intelligence (AI). There were significant differences by country ($x^{2}=84.89,df=9,p<0.001$). These responses were the highest across Australia (22.9%), the United States (18.3%), and Mexico (13.4%), and the lowest in France (2.9%) and Belgium (4.4%). References to the futuristic nature of deepfake technology, fear of its capabilities and accessibility, as well as the realness of deepfaked images, showed up across these responses. For example, one man from Mexico (aged 25-34) commented on how “it is difficult to assimilate [to] these types of tools, before they seemed futuristic but they are already very normal.” Many respondents described deepfakes as a troubling and terrifying advancement of AI technology. Some, particularly women, indicated their concern about being victimized by deepfake pornography. For example, a woman from Denmark (aged 25-34) referred to deepfakes as “widely scary” because “it’s hard not to think if you have become a victim yourself without knowing it.” Other respondents focused on the continued growth of deepfakes and the implications they might have for future generations. For example, an Australian woman (aged 45-54) described how she “fear[ed] for [her] daughter’s future.” Comments on the realness of deepfakes was a common thread throughout responses, with a Polish man (aged 35-44) saying “soon it will not be known if what we [are] seeing is reality or not.” Similarly, a Belgian woman (aged 45-54) questioned “can we still know if the photo or video seen or received is real or not?! It becomes difficult to make your own judgement.” Respondents also spoke to the difficulties of regulating deepfakes online. A man from Mexico (aged 65+) emphasized how it is “almost impossible to control,” while a woman from the United States (aged 65+) asked “how can one protect oneself from this? Or prove it’s not them?”
@@
The third most common theme was punitive and justice attitudes. Across responses, men were slightly more likely to express punitive and justice attitudes ($X^{2}=4.85,df=1,p<0.05$). References to punishment, injustice, criminalization, and governmental responsibility were common in South Korea. Just over half (55%) of South Korean respondents demanded strengthened laws relating to deepfakes. Responses often referred to deepfake pornography as a “sex crime” and called for the implementation of strong criminal justice regulations and sanctions. Accordingly, many South Korean respondents argued for the development of criminal sanctions that imposed “strong punishment” for perpetrators of deepfake pornography. As one South Korean respondent (male, aged 45-54) stated, “deepfakes [are] not yet controversial, but it can cause many social problems. In reality, legal regulations are needed to strongly punish [these] sex crimes.”
@@
Finally, only about 4% of responses indicated victim-blaming attitudes. These respondents noted that people should be more careful with what they post on the internet, and that they should refrain from taking or sharing nude photos or videos, and stop complaining about the issue and take responsibility. These responses were highest in Australia (32.0%) and lowest in South Korea (0.0%). For example, a non-binary respondent from Mexico (aged 35-44) and a woman from Poland (aged 55-64) noted that people have to be careful with what they share on the internet. Two Australian respondents encouraged people to stop taking nude images altogether; for example, one Australian man (aged 55-64) commented “if you don’t want something shared on the internet don’t take nude photos” while an Australian woman (aged 65+) said, “that’s why I do NOT send photos over the internet at all. I can’t understand why people send photos to anyone — keep your [images] private.” One man from Poland (aged 65+) was more explicit and indicated that people should not complain about being victimized: “you just should not share any photos on the web at all, and if you make it available, you should take into account the consequences and then regret pretending to be a fool and even worse to blame anyone.”
@@ 5. Discussion
Our study generated three main findings about NSII awareness, behaviors, and gender dynamics. First, despite widespread press coverage, particularly in Western countries, the concept of deepfake pornography is not universally known, yet on balance when informed about the concept, respondents thought behaviors associated with non-consensual fake intimate imagery should be criminalized. Second, the most common behaviors self-reported by respondents were passive and readily available (e.g., the consumption of celebrity deepfake pornography), with behaviors requiring effort and/or money being very rare (e.g., creating deepfake pornography). Third, compared to women, men think of these behaviors as less bad, and are more likely to also report being both victimized by and perpetrating NSII.
@@ 5.1. Attitudes and Awareness
Less than 30% of all respondents indicated some level of familiarity with the concept of deepfake pornography. This was also arguably reflected in the qualitative responses, with only 7.8% of all respondents answering the open-ended question. Presumably, as the technology becomes more widespread and ever more available, familiarity will increase. This may be a double-edged sword; it is likely to result in increased use of NSII as an abuse vector (both for purposes of humiliation, and as a tool for sexual extortion), but may also result in increased suspicion of the veracity of explicit content. One positive finding was general agreement across countries that perpetration behaviors associated with ”traditional” image-based sexual abuse are deserving of criminalization. Research has widely found some degree of victim-blaming reflected in attitudes towards traditional non-consensual intimate imagery (Flynn et al., 2023; Mortreux et al., 2019; Sugiura and Smith, 2020). Attitudes were slightly more lenient with regard to deepfake behaviors, although over time attitudes may evolve with increased familiarity. For instance, the realization that any private individual can be deepfaked, requiring no actions on the part of the target, may engender more empathy and less victim-blaming attitudes. It should be noted that few respondents in the qualitative part of the survey expressed victim-blaming attitudes. This can be compared to other more commonly expressed pro-social attitudes in the qualitative responses, including those expressing condemnation or support for punitive action.

### 13 Feb 2024 v2

@@ 4.2. Behaviors
Behaviors were asked about in three different ways. First, respondents who reported that they were quite familiar with the term “deepfakes” or knew a little bit about it (28.1%, $95\%CI=27.0\%-29.3\%$) were asked a follow-up question about whether they had engaged in specific deepfake pornography behaviors. Second, as part of the larger survey, respondents were asked about general image-based sexual abuse victimization and perpetration. If they reported experiences with digitally altered content, they were asked follow-up questions about whether the digitally altered content involved video(s), and if so, if the video(s) was a deepfake. This is distinct from the first question in that it a) asks for more details of the perpetration/victimization (e.g., whether they had created, threatened to share, or shared such content), and b) asks respondents who reported victimization to clarify whether they had been the victim of deepfake pornography specifically.
4.2.1. Behaviors of “Familiar” Respondents
The results of the first question (measuring perpetration behaviors) can be seen in Figure 2, where the reported data represents the percentage of all respondents. The respondents who did not see this question due to a lack of familiarity with deepfake pornography are represented in the missing data, which would make each bar sum to 100% if included. For example, 70% of respondents in Spain indicated they had never heard of this until now or they had only heard the phrase, so they were not asked this question. The most common behavior was the viewing of celebrity deepfake pornography, ranging from 2.8% ($95\%CI=2.1\%-3.7\%$) of all respondents in Poland, to 8.0% ($95\%CI=6.6\%-9.5\%$) of all respondents in Australia. Aggregated across the different countries, 6.2% ($95\%CI=5.6\%-6.9\%$) of all respondents indicated that they had viewed celebrity deepfake pornography. The second most common behavior (2.9% of all respondents, $95\%CI=2.5\%-3.3\%$) was the viewing of deepfake pornography involving “ordinary” people, ranging from 0.7% ($95\%CI=0.4\%-1.3\%$) in Denmark, to 5.1% ($95\%CI=4.1\%-6.2\%$) in Mexico. Behaviors that required more time, skills, or resources were rarer, such as creating or paying someone to create deepfake pornography of ordinary people (0.6%, $95\%CI=0.4\%-0.9\%$, and 0.5%, $95\%CI=0.3\%-0.7\%$), respectively, aggregated across countries)).
Figure 2. Deepfake Pornography Behaviors by Country. Percentage of all respondents who indicated engaging, or not, in a particular behavior. 95% confidence interval bars presented in grey.
@@ 4.2.1. Behaviors of “Familiar” Respondents
Figure 2. Deepfake Pornography Behaviors by Country. Percentage of all respondents who indicated engaging, or not, in a particular behavior. 95% confidence interval bars presented in grey.
@@ 4.3.1. Victimization
Experiences broken out by country are available in Figures 3 and 4. Aggregated across behaviors, respondents from Australia (3.7%), Mexico (2.9%), and South Korea (3.1%) reported the highest rates of victimization by deepfake pornography. Full results with 95% confidence intervals can be seen in Table A2 in the Appendix.
Aggregated across countries, 2.2% ($95\%CI=1.8\%-2.6\%$) of respondents reported some form of victimization. This included: 1.2% ($95\%CI=0.9\%-1.5\%$) of respondents who reported that someone had created deepfake pornography content of them; 1.3% ($95\%CI=1.0\%-1.6\%$) who reported that someone had posted or sent deepfake pornography content of them; and 1.2% ($95\%CI=0.9\%-1.5\%$) who reported that someone had threatened to post or send deepfake pornography content of them.
Figure 3. Percentage of Respondents Reporting Different Types of Deepfake Pornography Video Victimization, by Country. 95% CIs in grey.
@@
Figure 3. Percentage of Respondents Reporting Different Types of Deepfake Pornography Video Victimization, by Country. 95% CIs in grey.
@@ 4.3.2. Perpetration
Aggregating across behavior types (creating, posting, threatening to post), self-reported deepfake pornography perpetration was rare (1.8%, $95\%CI=1.4\%-2.2\%$), with the highest rates reported in Australia (2.4%) and the United States (2.6%), and lowest in Belgium ($0.3\%$). Full results can be seen in Table A2 in the Appendix. Aggregating across countries, 1.0% ($95\%CI=0.8\%-1.4\%$) of respondents indicated creating deepfake pornography, 1.0% ($95\%CI=0.8\%-1.3\%$) reported threatening to post, send, or share deepfake pornography, and 0.7% ($95\%CI=0.5\%-0.9\%$) reported actually posting, sending, or sharing deepfake pornography content. We included the creation of deepfake pornography as a perpetration behavior to mirror the fact that “creation” behaviors, such as filming or photographing real content without consent, fall under the umbrella of image-based sexual abuse perpetration. Nevertheless, we acknowledge that such behaviors might not be considered “perpetration” by others.
Figure 4. Percentage of Respondents Reporting Different Types of Deepfake Pornography Video Perpetration, by Country. 95% CIs in grey.
@@
Figure 4. Percentage of Respondents Reporting Different Types of Deepfake Pornography Video Perpetration, by Country. 95% CIs in grey.
@@ 4.4.1. Awareness and Attitudes
Given some existing literature on gendered attitudes and behaviors towards image-based sexual abuse more broadly, we wanted to assess whether gender played a role in respondents’ answers. There was more reported familiarity of deepfake pornography by respondents who identified as men, as opposed to those who identified as women. A Pearson’s chi-squared test with Rao & Scott adjustment confirmed the significant relationship between NSII awareness and gender ($F=58.7,ndf=3.0,ddf=49287.1,p<0.001)$. Men were 1.92 times ($2.4095\%CI=1.53-$ 2.40) more likely than women to report being “quite familiar with this” and were significantly less likely to report having never heard of deepfake pornography until now ($RR=0.69,95\%CI=0.65-0.73$). Full details can be seen in Figure A1 of the Appendix.
@@
Figure 5. Heatmap of Mean Criminalization Attitudes by Respondent Gender. Scale ranges from -2 (“Definitely should not be a crime”) to 2 (“Definitely should be a crime”), with a midpoint of 0 (“Not sure”). Brighter red is more deserving of criminalization, darker blue is less deserving of criminalization. Standard errors are presented in parentheses.
Heatmap showing attitudes towards various image-based sexual abuse behaviors. The x axis is gender (women and men), and the y axis is each type of behavior (e.g., posting, sending, or showing real intimate content without consent). The behaviors involving real content all fall between “should be a crime” and “definitely should be a crime,” and all weighted means are ¿0, indicating that all of the behaviors fell between “not sure” and “definitely should be a crime.”
Figure 5. Heatmap of Mean Criminalization Attitudes by Respondent Gender. Scale ranges from -2 (“Definitely should not be a crime”) to 2 (“Definitely should be a crime”), with a midpoint of 0 (“Not sure”). Brighter red is more deserving of criminalization, darker blue is less deserving of criminalization. Standard errors are presented in parentheses.
@@ 4.4.2. Behaviors
Of the two most commonly reported behaviors, men respondents were more likely to report viewing deepfake pornography; both that of celebrities ($F=71.1,ndf=1,ddf=16510,p<0.001,RR=3.5,95\%CI=2.77-4.42$) and non-celebrities ($F=14.3,ndf=1,ddf=16510,p<0.001,RR=2.66,95\%CI=1.91-3.70$).
Figure 6. Deepfake Pornography Behaviors by Gender. Percentage of all respondents who indicated engaging, or not, in a particular behavior. 95% confidence interval bars presented in grey.
@@
Figure 6. Deepfake Pornography Behaviors by Gender. Percentage of all respondents who indicated engaging, or not, in a particular behavior. 95% confidence interval bars presented in grey.
@@ 4.4.3. Respondents Who Reported Personal Victimization or Perpetration
Looking at the effects of gender on victimization and perpetration, men reported significantly higher rates across all victimization and perpetration experiences related to creation and threats. For example, the risk of men reporting having been threatened with deepfake dissemination was 2.6 times that of women, and the risk of them threatening someone else was 1.9 times that of women. Full results are presented in Table 4.4.3. Notably, there were no significant differences by gender in the risk of disseminating content or having their content disseminated.
@@ 4.5. Qualitative Responses
In the survey, respondents were asked an optional, open-ended question: “Would you like to add any additional thoughts on the topic of ‘deepfakes’? Please feel free to skip this question if you don’t have anything to add.” A total of 2,477 open-ended responses were collected. Of those responses, 1,180 were removed due to inappropriate or irrelevant responses. This left us with 1,296 open-ended responses (7.8% of respondents). No respondents discussed their own experiences of either victimization or perpetration.
@@
The second most prevalent theme, found in just over 30% of all responses to the open-ended question, referred to a fear of technology; specifically deepfakes and artificial intelligence (AI). There were significant differences by country ($x^{2}=32.81,df=9,p<0.001$), but not by gender ($x^{2}=0.011,df=1,p=0.73$). These responses were the highest in the Netherlands (50.0%), and the lowest in France (16.2%) and South Korea (19.7%). References to the futuristic nature of deepfake technology, fear of its capabilities and accessibility, as well as the realness of deepfaked images, showed up across these responses. For example, one man from Mexico (aged 25-34) commented on how “it is difficult to assimilate [to] these types of tools, before they seemed futuristic but they are already very normal.” Many respondents described deepfakes as a troubling and terrifying advancement of AI technology. Some, particularly women, indicated their concern about being victimized by deepfake pornography. For example, a woman from Denmark (aged 25-34) referred to deepfakes as “widely scary” because “it’s hard not to think if you have become a victim yourself without knowing it.” Other respondents focused on the continued growth of deepfakes and the implications they might have for future generations. For example, an Australian woman (aged 45-54) described how she “fear[ed] for [her] daughter’s future.” Comments on the realness of deepfakes was a common thread throughout responses, with a Polish man (aged 35-44) saying “soon it will not be known if what we [are] seeing is reality or not.” Similarly, a Belgian woman (aged 45-54) questioned “can we still know if the photo or video seen or received is real or not?! It becomes difficult to make your own judgement.” Respondents also spoke to the difficulties of regulating deepfakes online. A man from Mexico (aged 65+) emphasized how it is “almost impossible to control,” while a woman from the United States (aged 65+) asked “how can one protect oneself from this? Or prove it’s not them?”
@@
The third most common theme was punitive and justice attitudes, seen in just over a quarter of respondents. Across responses, men were slightly more likely to express punitive and justice attitudes ($X^{2}=4.85,df=1,p<0.05$), and there were significant differences by country ($X^{2}=91.18,df=9,p<0.001$). References to punishment, injustice, criminalization, and governmental responsibility were common in South Korea. Just over half (55.3%) of South Korean respondents demanded strengthened laws relating to deepfakes. Responses often referred to deepfake pornography as a “sex crime” and called for the implementation of strong criminal justice regulations and sanctions. Accordingly, many South Korean respondents argued for the development of criminal sanctions that imposed “strong punishment”” for perpetrators of deepfake pornography. As one South Korean respondent (male, aged 45-54) stated, “deepfakes [are] not yet controversial, but it can cause many social problems. In reality, legal regulations are needed to strongly punish [these] sex crimes.”
@@
Finally, only about 4% of responses indicated victim-blaming attitudes. These respondents noted that people should be more careful with what they post on the internet, and that they should refrain from taking or sharing nude photos or videos, and stop complaining about the issue and take responsibility. There were no significant differences by gender ($X^{2}=1.68,df=1,p=0.20$), nor by country (per a Fisher’s exact test for count data, $p=0.08$). For example, a non-binary respondent from Mexico (aged 35-44) and a woman from Poland (aged 55-64) noted that people have to be careful with what they share on the internet. Two Australian respondents encouraged people to stop taking nude images altogether; for example, one Australian man (aged 55-64) commented “if you don’t want something shared on the internet don’t take nude photos” while an Australian woman (aged 65+) said, “that’s why I do NOT send photos over the internet at all. I can’t understand why people send photos to anyone — keep your [images] private.” One man from Poland (aged 65+) was more explicit and indicated that people should not complain about being victimized: “you just should not share any photos on the web at all, and if you make it available, you should take into account the consequences and then regret pretending to be a fool and even worse to blame anyone.”
@@ 5. Discussion
Our study generated four main findings about AI-IBSA/deepfake pornography awareness, behaviors, and gender dynamics. First, despite widespread press coverage, particularly in Western countries, the concept of deepfake pornography is not well-known. Yet on balance, when informed about the concept, respondents thought behaviors associated with non-consensual fake intimate imagery should be criminalized. Second, self-reported victimization prevalence was relatively low, as compared to more traditional image-based sexual abuse prevalence rates. Third, the most common behaviors self-reported by respondents were passive and readily available (e.g., the consumption of celebrity deepfake pornography), with behaviors requiring effort and/or money being very rare (e.g., creating deepfake pornography). Finally, compared to women, men think of these behaviors as less bad, and are also more likely to report being both victimized by and perpetrating AI-IBSA.
@@ 5.1. Attitudes and Awareness
Less than 30% of all respondents indicated some level of familiarity with the concept of deepfake pornography. This was also arguably reflected in the qualitative responses, with only 7.8% of all respondents answering the open-ended question. Presumably, as the technology becomes more widespread and ever more available, familiarity will increase. This may be a double-edged sword; it is likely to result in increased use of AI-IBSA as an abuse vector (both for purposes of humiliation, and as a tool for sexual extortion), but may also result in increased suspicion of the veracity of explicit content. One positive finding was general agreement across countries that perpetration behaviors associated with “traditional” image-based sexual abuse are deserving of criminalization. Research has widely found some degree of victim-blaming reflected in attitudes towards traditional non-consensual intimate imagery (Flynn et al., 2023; Mortreux et al., 2019; Sugiura and Smith, 2020). Attitudes were slightly more lenient with regard to deepfake behaviors, although over time attitudes may evolve with increased familiarity. For instance, the realization that any private individual can be deepfaked, requiring no actions on the part of the target, may engender more empathy and less victim-blaming attitudes. It should be noted that few respondents in the qualitative part of the survey expressed victim-blaming attitudes. This can be compared to other more commonly expressed pro-social attitudes in the qualitative responses, including those expressing condemnation or support for punitive action.